1. A certain company makes electrical cables having a mean strength of 40 kg/cm² and a standard deviation of 2 kg/cm². Assuming the strength follows a normal distribution,




a. What percentage of cables will have a strength exceeding 43 kg/cm²?




b. If strength is measured to the nearest kg/cm², what percentage of cables will have




strength exceeding 43 kg/cm²?




c. Illustrate the normal curve for a and b.

1
Expert's answer
2022-06-13T09:44:05-0400

a)


"P(X>43)=1-P(Z\le\dfrac{43-40}{2})"

"=1-P(Z\le1.5)=0.0668"

b)


"P(X\ge43.5)=1-P(Z<\dfrac{43.5-40}{2})"

"=1-P(Z<1.75)=0.0401"


c)

a)



b)